OpenStudy (clamin):

PLEASE HELP!! MEDAL!! What is the solution to the inequality below?? x^2 > 100 a) -10 less than or equal to x less than or equal to 10 b) x<-10 or x>10 c)-10

OpenStudy (haseeb96):

x^2>100 taking square root on both sides then the above equation would be x > +10 x < -10 so the correct answer would be ????

OpenStudy (michele_laino):

hint: \[{x^2} - 100 = \left( {x - 10} \right)\left( {x + 10} \right)\]

OpenStudy (clamin):

b